OpenStudy (anonymous):

Use the quadratic formula to solve for x. 3x^2+6x=-2

OpenStudy (anonymous):

@jim_thompson5910 @satellite73 @tejasvir @TheSmartOne @Walkerman15

jimthompson5910 (jim_thompson5910):

3x^2+6x=-2 3x^2+6x+2=-2+2 3x^2+6x+2=0 use the quadratic formula \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] to solve for x. In this case, a = 3, b = 6, c = 2